OpenStudy (anonymous):

Solve each system of equations with substitution -6a-5v=46 -6a-4v=44

OpenStudy (anonymous):

You can solve for -6a by adding 4v to both sides of equation 2, giving you\[-6a=44+4v\] Plug this -6a into the first equation to give you \[(44+4v)-5v=46\] Solve for v \[v=-2\] plug new v into the original second equation: \[-6a-4(-2)=44\]solving for a gives you\[a=-6\]